# Addressing the Non-perturbative Regime of the Quantum Anharmonic Oscillator by Physics-Informed Neural Networks

### v1

@@ Abstract
The use of deep learning in physical sciences has recently boosted the ability of researcher to tackle physical systems where little or no analytical insight is available. Recently, the Physics--Informed Neural Networks (PINNs) have been introduced as one of the most promising tools to solve systems of differential equations guided by some physically grounded constraints. In the quantum realm, such approach paves the way to a novel approach to solve the Schrödinger equation for non-integrable systems. By following an unsupervised learning approach, we apply the PINNs to the anharmonic oscillator in which an interaction term proportional to the fourth power of the position coordinate is present. We compute the eigenenergies and the corresponding eigenfunctions while varying the weight of the quartic interaction. We bridge our solutions to the regime where both the perturbative and the strong coupling theory work, including the pure quartic oscillator. We investigate systems with real and imaginary frequency, laying the foundation for novel numerical methods to tackle problems emerging in quantum field theory.
@@ I Introduction
Much of the modern understanding in physics is due to the theoretical framework posed in place by quantum field theories. Ranging from the quantization of the electromagnetic field to the explanation of heat conduction in solids, the ingenious and revolutionary trick was to rethink the fields as a collection of harmonic oscillators [1]. Since harmonic systems are one of the few integrable systems either in classical or quantum mechanics, and even if the solution of harmonic systems is an academic exercise, it is also wide knowledge how predictive these systems are. This implies that analytical computations allow the knowledge of their entire spectrum and to gain insight from a multitude of physical systems thanks to the mapping between bosonic systems and the quantum harmonic oscillator [2]. This boosted not only the acclaimed achievements in the quantum theory of radiation [3, 4], but also the application of field theoretical tools to treat the small oscillations in molecules [5] or in the solid lattice and fostered our understanding of condensed matter. [6].
@@
However, if the harmonic potential is modified by the addition of terms proportional to higher powers of the position coordinates, the system loses almost always its integrability  [^1] It is worth noting that the Morse potential (and several of its variations) can account for the anharmonic behavior and are still integrable. [7]. Nevertheless, since the pioneering work by lord Rayleigh in classical physics [8] and subsequently by Schrödinger [9] in the quantum context, it was shown that slight perturbation breaking the harmonicity can be successfully treated within the framework of perturbation theory. In the present paper, we aim to provide a deep-learning perspective on a system that has long served as a prototypical interaction model [10, 11], and it served and still serves to study innovative techniques in quantum field theory [12, 13, 14, 15, 16]. The anharmonic oscillator with a term proportional to $x^{4}$ can be indeed seen as a $\lambda\phi^{4}$ theory in $(0+1)-$dimensions. Furthermore, such a system has been consistently used in several contexts, most notably in solid-state physics to account for the hardening of the phonon dispersion relation [17, 18, 19], to investigate the formation of molecules with hydrogen bonds and their vibrational modes [20, 21], but also for foundational studies such as nonlinear quantum mechanics [22, 23], quantum chaos and control [24, 25].
@@
Some of us have already developed both supervised [26, 27, 28] and unsupervised [29] machine learning methods, including supervised quantum machine learning to address ground state classification [30, 31, 32] of a physical system. Here, we aim to address the unsupervised solution of the Schrödinger [33] equation, to explore the potential of Physics-Informed Neural Networks (PINNs) for solving Partial Differential Equations (PDEs) as a paramount representative application to nontrivial physical phenomena.
@@
We applied the method to find both the ground state and the excited states of the quantum harmonic and anharmonic oscillator, with real or imaginary frequency. We employed a PINN that splits into two separate networks [34], one for the eigenvalue and one for the eigenfunction. Furthermore, it also employs an auxiliary output [35] to obtain the normalization integral. We also employed transfer learning, utilizing the weights and biases of the already trained networks to obtain an excellent initialization for networks with similar parameters.
@@
The remaining part of the paper is organized as follows. In Section II, we briefly introduce the details of the deep learning model employed, explaining in particular loss functions and architecture of the neural network. Section III gives the detail of the quantum anharmonic oscillator, and delves into the detailed explanation of the unsupervised learning strategy that we followed. Section IV shows the results for the unperturbed (harmonic) oscillator while in Section V the results for the anharmonic oscillator are described. Section VI shows the results for the pure quartic oscillator while in Section VII we address the scaling behavior of the eigenenergies in function of $\lambda$. Lastly, Section VIII concludes the paper and address some interesting perspectives for future research directions.
@@ II.1 Physics-informed neural networks
$f=MSE_{PDE}+MSE_{data}+MSE_{phys}$ (1)
@@
Where $MSE_{data}$ is the standard loss for a Neural Network  [^2] In our case this loss will not be present since we are working in an unsupervised setting. $MSE_{phys}$ corresponds to the losses given by the physical constraints of the system, like boundary conditions and initial conditions. Lastly, $MSE_{PDE}$ is the mean squared error of the differential equation. For instance, given a PDE with implicit solution $\mu(t)$ that depends on $\mu(t)$ and its derivative $\partial_{t}\mu(t)$ with respect to the independent variable $t$, the differential equation can be written as:
$\partial_{t}\mu(t)+N[\mu;\lambda]=0$ (2)
@@
where N is a nonlinear operator and $\lambda$ is some set of parameters. Given this, we can write $MSE_{PDE}$ for a network designed to solve this equation as:
$\frac{1}{n}\sum_{i=1}^{n}(\partial_{t}\mu(t_{i})+N[\mu(t_{i});\lambda])^{2}$ (3)
@@
Where the ${t_{i}}$ is the set of n points, called collocation points, where the PDE will be evaluated. This makes it possible to train neural networks even in the presence of noisy or very little data, and even in a fully unsupervised setting, such as in this paper. These networks have had a number of applications in various fields. Just to name a few applications of PINNs, the most recent and relevant use cases are in fluid dynamics [37, 38, 39], heat transfer problems [40], scattering in composite optical materials [41], oscillating dynamics of black holes ([42]), reconstruction of top quark kinematic properties ([43]) and to compute the electronic density in complex molecular systems ([44]).
II.2 The quantum anharmonic oscillator
@@
The anharmonic oscillator model has played a pivotal role in many research areas in physics, mostly due to its simplicity but especially because it is a nontrivial nonlinear problem. It appeared to be crucial in the development of several branches of physics ranging from high-energy particle physics to low-energy molecular dynamics. The Hamiltonian describing the problem in one-dimension for a system of unitary mass is:
@@
in which $\omega$ is the harmonic frequency and $\lambda$ quantifies the strength of the interaction-like term, and finally $p$ and $x$ are the standard position and momentum operator satisfying $[x,p]=i$. The potential $V(x)$ could assume two relevant shapes for physical applications, depending on the signs of $\omega^{2}$ and $\omega^{2}/\lambda$. When both terms are positive the potential has only one minimum and its shape is that of a “super parabola”, whereas when the frequency is imaginary, we are in the presence of a double-well potential that is the archetype of a model exhibiting a spontaneously symmetry-breaking mechanism. Without losing generality in the following we will assume that the oscillator has unitary mass and we will work in natural units $m=c=\hbar=1$, where $m$ is the mass of the oscillator, $c$ and $\hbar$ are the speed of light in the vacuum and the reduced Planck constant respectively.
@@ III Methodology
To obtain eigenfunctions and eigenstates for the anharmonic oscillator, we employ two neural networks that will be trained in parallel. Both networks take the same input, the position $x$, which will be sampled in an interval $[-\frac{L}{2},\frac{L}{2}]$ where $L$ is a hyperparameter that controls the size of the computational domain and has to be chosen large enough that the wave function is vanishingly small at the system’s boundary, but not so large that the network has to predict a vast area of just zeros.
The first neural network will output the wave function $\psi(x)$ and the auxiliary output $\nu(x)$, which is defined as:
@@
and will be used to ensure normalization. On the other hand, the second network will output a vector of guesses for the energy. We take as the actual guess for the eigenvalue the average of the outputs of the second Neural Network.
We need to optimize a total of 7 losses to train the network. First of all, we need a way to compute the integral of the square module of the eigenfunction to be able to normalize the wave function. To do it in a mesh-free way we utilize an auxiliary output, $\nu(x)$, and train the network to make it a guess for integral from 00 to $x$ of the square module of the outputted wave function, (see equation (5)). To ensure that this auxiliary output has the desired form, we utilize automatic differentiation to obtain the partial derivative of the auxiliary output, and require it to be equal to the squared modulus of the wave function, thus obtaining the integral loss:
@@
where $\psi_{i}(x)$ are the known eigenfunctions. This loss will always be 0 when training the first network, the one that will be trained to output the ground state, but each time a network converges it will be saved and utilized to obtain the $\psi_{i}(x)$ utilized to compute this loss in the training of the networks of higher energy states. At this point, we need to include the loss that will allow the network to gauge the correctness of the outputted eigenstate and eigenvalue for the current potential, that is the differential equation loss:
@@
$e^{a(E_{PINN}-E_{init})}=0,$ (11)
where $E_{PINN}$ is the predicted energy, $a$ is a hyperparameter, and $E_{init}$ is either another hyperparameter, when $n=0$, or the energy of the previous state, for excited states. The choice to increase $E_{init}$ as the quantum number increases has been taken to avoid an excessive weight for this loss for a higher energy state, which might cause the network to fail to converge at all. The energy minimization loss, however, will be greater than 00 even when we reach the correct solution, and that will cause the network to converge to an incorrect state. In order to avoid this we gradually decrease this loss’ weight as the training continues. In this way in the earlier epochs, this loss will steer the network towards lower energy states, while for the latter epochs it will become negligible and allow the network to settle to the correct solution.
The last loss that we implement is one that is not actually needed to obtain a solution, but is instead added to allow the network to converge much faster, acting as an inductive bias: the symmetry loss. Since the potential is symmetric with respect to 0, the wave functions that solve Eq. (10) will also have to be alternatively symmetric and anti-symmetric with respect to 0. We can enforce this condition by adding the loss:
@@
For the architecture, the network that outputs the wave function and the integral is made up of $7$ hidden layers of $256$ neurons each while the one used to predict the energy is made up of $4$ layers, also of $256$ neurons. Batches of $512$ collocation points were employed. We want each batch to span the whole domain, but we also want to utilize fully the mesh-free nature of the neural method. To do so we construct a grid of $512$ equally spaced points, and for each of them, we sample randomly in an interval centered on it, thus obtaining a batch that spans the whole domain but whose values change slightly at each iteration, ideal characteristics to optimize the network’s convergence.
@@ III.1 Weights
The Orthogonality loss is proportional to the quantum number of the excited state we want to study.
@@ III.2 Evaluation metrics
where $E_{Ex}$ is the analytic value of the energy, while $E_{PINN}$ is the one predicted by the network. On the other hand to evaluate the eigenvector the fidelity between the predicted wave function and the exact one [50, 51, 52] is employed:
@@
where the $x_{i}$ are a set of points spanning the whole domain, and $\psi_{Ex}(x)$ and $\psi_{PINN}(x)$ are the values of the wave function at the point $x$ as given by the analytic expression and by the network, respectively. Both of the vectors will be normalized before calculating the similarity, to avoid getting incorrect results due to inaccurate scaling. This should equal one when the two wave functions are the same. These metrics, however, might depend on the choice of evaluation points. Therefore once the training has converged we compute them $1000$ times for different sets of points, sampled in the same way we sample the training batches, to make sure that we always get a different set of points, and that each of these sets still spans the whole domain, to obtain a valid estimate of the fidelity. The final values will be the average of these estimates and will come with a standard deviation quantifying the dependence on the choice of evaluation points. For a deeper analysis, in which we discuss better the variance of these results we refer the reader to Section VII.

### v2

@@ Abstract
The use of deep learning in physical sciences has recently boosted the ability of researchers to tackle physical systems where little or no analytical insight is available. Recently, the Physics--Informed Neural Networks (PINNs) have been introduced as one of the most promising tools to solve systems of differential equations guided by some physically grounded constraints. In the quantum realm, such an approach paves the way to a novel approach to solve the Schrödinger equation for non-integrable systems. By following an unsupervised learning approach, we apply the PINNs to the anharmonic oscillator in which an interaction term proportional to the fourth power of the position coordinate is present. We compute the eigenenergies and the corresponding eigenfunctions while varying the weight of the quartic interaction. We bridge our solutions to the regime where both the perturbative and the strong coupling theory work, including the pure quartic oscillator. We investigate systems with real and imaginary frequency, laying the foundation for novel numerical methods to tackle problems emerging in quantum field theory.
@@ I Introduction
Much of the modern understanding in physics is due to the theoretical framework established by by quantum field theories. Ranging from the quantization of the electromagnetic field to the explanation of heat conduction in solids, the ingenious and revolutionary trick was to rethink the fields as a collection of harmonic oscillators [1]. Since harmonic systems are one of the few integrable systems either in classical or quantum mechanics, and even if the solution of harmonic systems is an academic exercise, it is also wide knowledge how predictive these systems are. This implies that analytical computations allow the knowledge of their entire spectrum and to gain insight from a multitude of physical systems thanks to the mapping between bosonic systems and the quantum harmonic oscillator [2]. This boosted not only the acclaimed achievements in the quantum theory of radiation [3, 4], but also the application of field theoretical tools to treat the small oscillations in molecules [5] or in the solid lattice and fostered our understanding of condensed matter. [6].
@@
However, if the harmonic potential is modified by the addition of terms proportional to higher powers of the position coordinates, the system almost always loses its integrability  [^1] It is worth noting that the Morse potential (and several of its variations) can account for the anharmonic behavior and are still integrable. [7]. Nevertheless, since the pioneering work by Lord Rayleigh in classical physics [8] and subsequently by Schrödinger [9] in the quantum context, it was shown that slight perturbation breaking the harmonicity can be successfully treated within the framework of perturbation theory. In the present paper, we aim to provide a deep-learning perspective on a system that has long served as a prototypical interaction model [10, 11], and it served and still serves to study innovative techniques in quantum field theory [12, 13, 14, 15, 16]. The anharmonic oscillator with a term proportional to $x^{4}$ can be indeed seen as a $\lambda\phi^{4}$ theory in $(0+1)-$dimensions. Furthermore, such a system has been consistently used in several contexts, most notably in solid-state physics to account for the hardening of the phonon dispersion relation [17, 18, 19], to investigate the formation of molecules with hydrogen bonds and their vibrational modes [20, 21], and also for foundational studies such as nonlinear quantum mechanics [22, 23], quantum chaos and control [24, 25].
@@
Some of us have already developed both supervised [26, 27, 28] and unsupervised [29] machine learning methods, including supervised quantum machine learning to address ground state classification [30, 31, 32] of a physical system. Here, we aim to address the unsupervised solution of the Schrödinger [33] equation, to explore the potential of Physics-Informed Neural Networks (PINNs) for solving Partial Differential Equations (PDEs) as a paramount representative application to nontrivial physical phenomena [34].
@@
We applied the method to find both the ground state and the excited states of the quantum harmonic and anharmonic oscillator, with real or imaginary frequency. We employed a PINN that splits into two separate networks [35], one for the eigenvalue and one for the eigenfunction. Furthermore, it also employs an auxiliary output [36] to obtain the normalization integral. We also employed transfer learning, utilizing the weights and biases of the already trained networks to obtain an excellent initialization for networks with similar parameters of the system Hamiltonian.
@@
The remaining part of the paper is organized as follows. In Section II, we briefly introduce the details of the deep learning model employed, explaining in particular loss functions and architecture of the neural network. Section III gives the detail of the quantum anharmonic oscillator, and delves into the detailed explanation of the unsupervised learning strategy that we followed. Section IV shows the results for the unperturbed (harmonic) oscillator while in Section V the results for the anharmonic oscillator are described. Section VI shows the results for the pure quartic oscillator while in Section VII we address the scaling behavior of the eigenenergies as a function of $\lambda$, the interaction strength. Lastly, Section VIII concludes the paper and address some interesting perspectives for future research directions.
@@ II.1 Physics-informed neural networks
$f={\mathcal{L}}_{PDE}+{\mathcal{L}}_{data}+{\mathcal{L}}_{phys}$ (1)
@@
Where the term ${\mathcal{L}}_{data}$ is the standard loss for a Neural Network  [^2] In our case this loss will not be present since we are working in an unsupervised setting.. ${\mathcal{L}}_{phys}$ corresponds to the losses given by the physical constraints of the system, like boundary conditions and initial conditions. Lastly, ${\mathcal{L}}_{PDE}$ is the mean squared error of the differential equation. For instance, given a PDE with implicit solution $\mu(t)$ that depends on $\mu(t)$ and its derivative $\partial_{t}\mu(t)$ with respect to the independent variable $t$, the differential equation can be written as:
$\partial_{t}\mu(t)+N[\mu;\xi]=0$ (2)
@@
where $N$ is a nonlinear operator and $\xi$ is some set of parameters. Given this, we can write ${\mathcal{L}}_{PDE}$ for a network designed to solve this equation as:
$\frac{1}{n}\sum_{i=1}^{n}(\partial_{t}{\mu_{net}}(t_{i})+N[{\mu_{net}}(t_{i});\xi])^{2}$ (3)
@@
Where the ${t_{i}}$ is the set of $n$ points, called collocation points, where the PDE will be evaluated. This makes it possible to train neural networks even in the presence of noisy or very little data, and even in a fully unsupervised setting, such as in this paper. These networks have had a number of applications in various fields. Just to name a few applications of PINNs, the most recent and relevant use cases are in fluid dynamics [37, 38, 39], heat transfer problems [40], scattering in composite optical materials [41], oscillating dynamics of black holes [42], reconstruction of top quark kinematic properties [43] and to compute the electronic density in complex molecular systems [44].
II.2 Related works
The Schrödinger equation stands at the cornerstone of the modern quantum theory and its subsequent development that led to the first and second quantum revolutions. Alas, the cases in which it is analytically solvable are only a few, and the central potential problem stands among them. This boosted the necessity to look for alternative ways of solving the differential equation giving the dynamics of a quantum object. In recent years, the enthusiasm around the potential of artificial intelligence methods triggered their applications also in quantum science and technology (see [45, 46] for thorough discussions). Several attempts have already been made to address the Schrödinger equation via Artificial Neural Networks. For instance, the authors in [47] already had some success in solving the dynamics of a system under the action of the Morse potential, anharmonic but still integrable. When moving to the particular class of neural networks considered here, unsupervised PINNs have already proven to be successful in addressing simple systems [33] as well as the Schrödinger equation with random potentials [35]. Moreover, PINNs have also been used to study the Schrödinger equation with time-dependent potentials [48] and to obtain solutions of the nonlinear one that arises in wave-guided quantum optics [49, 50, 51]. Deep complex neural networks [52] are also of special interest for future applications, due to their ability to naturally address complex-valued wavefunctions. They also have been used to discover the governing PDE of a system from data [53].
II.3 The quantum anharmonic oscillator
@@
The anharmonic oscillator model has played a pivotal role in many research areas in physics, mostly due to its simplicity but especially because it is a nontrivial nonlinear problem. It appeared to be crucial in the development of several branches of physics ranging from high-energy particle physics to low-energy molecular dynamics.
Before delving into the discussion of our main problem, we review briefly the terminology. The solution of the Schrödinger equation provides the possible energy levels and the corresponding wave functions, describing the dynamics of a quantum object in a given potential field. The time-independent Schrödinger Equation allows to find the the stationary states of a quantum system, namely those for which the energy remains constant over time. To write it down, one needs to introduce the Hamiltonian operator $\hat{H}$ that represents the total energy of a quantum system, such that:
$\hat{H}\psi(\mathbf{r})=E\psi(\mathbf{r})$ (4)
$\psi(\mathbf{r})$ is the wave function that describes the spatial distribution of the particle, and $E$ is the energy eigenvalue associated with that state.
[…]
$\hat{H}=\hat{K}+\hat{V}=-\frac{\hbar^{2}}{2m}\nabla^{2}+V(\mathbf{r}),$
where we have substituted the momentum operator $\hat{\vec{p}}$ with the corresponding a differential operator in the position representation, i.e., $\hat{\vec{p}}\rightarrow-i\hbar\nabla$ From now on for the sake of simplicity, we will omit the hat symbol on the operators.
The Hamiltonian accounting for the total energy of the anharmonic oscillator in one-dimension for a system of unitary mass is:
@@
in which $\omega$ is the harmonic frequency and $\lambda$ quantifies the strength of the interaction-like term, and finally $p$ and $x$ are the standard position and momentum operators in one dimension satisfying $[x,p]=i$. The potential $V(x)$ could assume two relevant shapes for physical applications, depending on the signs of $\omega^{2}$ and $\omega^{2}/\lambda$. When both terms are positive the potential has only one minimum and its shape is that of a “super parabola”, whereas when the frequency is imaginary, we are in the presence of a double-well potential that is the archetype of a model exhibiting a spontaneously symmetry-breaking mechanism. Without losing generality in the following we will assume that the oscillator has unitary mass and we will work in natural units $m=c=\hbar=1$, where $m$ is the mass of the oscillator, $c$ and $\hbar$ are the speed of light in the vacuum and the reduced Planck constant respectively.
@@ III Methodology
To obtain eigenfunctions and eigenstates for the anharmonic oscillator, we employ two neural networks that will be trained in parallel. The first and main network takes as input the position $x$, which will be sampled in an interval $[-\frac{L}{2},\frac{L}{2}]$ where $L$ is a hyperparameter that controls the size of the computational domain and has to be chosen large enough that the wave function is vanishingly small at the system’s boundary, but not so large that the network has to predict a vast area of just zeros.
It will then output the wave function $\psi(x)$ and the auxiliary output $\nu(x)$, which is defined as:
@@
and will be used to ensure normalization. On the other hand, the second network takes as input just the constant value $1$ and outputs a guess for the energy.
We need to optimize a total of seven losses to train the networks. First of all, we need a way to compute the integral of the squared modulus of the eigenfunction to be able to normalize the wave function. To do it in a mesh-free way we utilize an auxiliary output, $\nu(x)$, and train the network to make it a guess for integral from zero to $x$ of the squared modulus of the outputted wave function, (see equation (6)). To ensure that this auxiliary output has the desired form, we utilize automatic differentiation to obtain the partial derivative of the auxiliary output, and require it to be equal to the squared modulus of the wave function, thus obtaining the integral loss:
@@
where $\psi_{i}(x)$ are the known eigenfunctions. For the case of the ground state the orthogonality loss will always be zero. This is because there are no discovered states to compare the first output to. Each time a network converges it will be saved and utilized to obtain the $\psi_{i}(x)$ utilized to compute this loss in the training of the networks of higher energy states. At this point, we need to include the loss that will allow the network to gauge the correctness of the outputted eigenstate and eigenvalue for the current potential, that is the differential equation loss:
@@
${e^{a(E_{PINN}-E_{init})}}$ (12)
where $E_{PINN}$ is the predicted energy, $a$ is a hyperparameter, and $E_{init}$ is either another hyperparameter, when $n=0$, or the energy of the previous state, for excited states. The choice to increase $E_{init}$ as the quantum number increases has been taken to avoid an excessive weight for this loss for a higher energy state, which might cause the network to fail to converge at all. We employ the exponential of the difference of the $E_{PINN}$ and $E_{init}$ because it is a monotonically increasing positive function in [,+][$\infty$,$\infty$][ - ∞ , + ∞ ]. This guarantees a loss that is both always positive and always pushes the network to reduce the energy. Thus, it should lead the network to discover the minimal energy state. The energy minimization loss, however, will be greater than zero even when we reach the correct solution, and that will cause the network to converge to an incorrect state. In order to avoid this we gradually decrease this loss’ weight as the training continues. In this way in the earlier epochs, this loss will steer the network towards lower energy states, while for the latter epochs it will become negligible and allow the network to settle to the correct solution.
The last loss that we implement is one that is not actually needed to obtain a solution, but is instead added to allow the network to converge much faster, acting as an inductive bias: the symmetry loss. Since the potential is symmetric with respect to 0, the wave functions that solve Eq. (11) will also have to be alternatively symmetric and anti-symmetric with respect to 0. We can enforce this condition by adding the condition:
@@
For the architecture, the main network, that is, the one that outputs the wave function and the integral, is made up of seven hidden layers of $256$ neurons each. The one used to predict the energy, on the other hand, is made up of four hidden layers also of $256$ neurons each. Even if this network only has to perform a linear transformation, empirically the presence of hidden layers seems to vastly improve convergence. Batches of $512$ collocation points were employed. We want each batch to span the whole domain, but we also want to fully utilize the mesh-free nature of the neural method. To do so we construct a grid of $512$ equally spaced points, and for each of them, we sample randomly in an interval centered on it, thus obtaining a batch that spans the whole domain but whose values change slightly at each iteration, ideal characteristics to optimize the network’s convergence. For the activation function, we employed the hyperbolic tangent, which is an infinitely differentiable function. Utilizing such an activation function is needed because otherwise the network would be unable to compute higher-order derivatives via automatic differentiation, which would prevent it from properly computing the loss in Eq. (11). As for the optimizer, we utilized ADAM (ADaptive Moment Estimation) [59]. Lastly we want to address the metrics used to compute the losses. It is not necessary to compute all the partial losses using the same metric. Instead, we utilized Sum of Absolute Errors for the losses enforcing the physical constraints of the system: the boundary conditions loss and the normalization loss. On the other hand we utilized Mean Squared Error for the integral, differential equation, orthogonality and symmetry losses. The energy minimization loss is just computed with Eq. (12). We choose to utilize SAE instead of the more standard MSE for the constraints in order to avoid the loss vanishing two quickly as the neural network outputs get closer to the desired output. This ensures that these conditions are fulfilled, making the rest of the training valid. This architecture is summarized in Figure 1.
@@ III.1 Weights
The Orthogonality loss is proportional to the quantum number of the excited state we want to study. See also [60] for a similar approach.
@@ III.2 Evaluation metrics
where $E_{Ex}$ is the analytic value of the energy, while $E_{PINN}$ is the one predicted by the network [^3] Here we do not use a norm and retain the sign because it provides additional information about the nature of the error. Namely, a positive $err_{E}$ means that the PINN underestimates the energy, while a negative $err_{E}$ signals an overestimation of the energy.. On the other hand to evaluate the eigenvector the fidelity between the predicted wave function and the exact one [61, 62, 63] is employed:
@@
where the $x_{i}$ are a set of points on an evenly spaced grid spanning the whole domain, and $\psi_{Ex}(x)$ and $\psi_{PINN}(x)$ are the values of the wave function at the point $x$ as given by the analytic expression and by the network, respectively. Both of the vectors will be normalized before calculating the similarity, to avoid getting incorrect results due to inaccurate scaling. This should equal one when the two wave functions are the same. These metrics, however, might depend on the choice of evaluation points. Therefore once the training has converged we compute them $1000$ times for different sets of points, sampled in the same way we sample the training batches, to make sure that we always get a different set of points, and that each of these sets still spans the whole domain, to obtain a valid estimate of the fidelity. The final values will be the average of these estimates and will come with a standard deviation quantifying the dependence on the choice of evaluation points. For a deeper analysis, in which we discuss better the variance of these results we refer the reader to Section VII.